how many liters of 60 percent acid solution has to be added to a 75 percent acid solution to make 20 liters of 72 percent?
Answer by mananth(16949)

percent ---------------- quantity
Acid type I 60 ----------------x liters
Acid type II 75 ----------------20-x liters
Mixture 72.00% ----------------20

60x+75( 20-x)=72*20
60x+1500-75x=1440
60x-75x =1440-1500
-15x= -60
/-15
x=4 liters 60.00% Acid type I
16 liters 75.00% Acid type II
m.ananth@hotmail.ca